OpenStudy (anonymous):

14x - 6 = -3 + 2x + 5

OpenStudy (anonymous):

looong decimal, but cancel out numbers to start with

OpenStudy (anonymous):

12x - 6 = -3 + 5

OpenStudy (anonymous):

then from there, the equation becomes 12x=6-3+5 by adding 6 to both sides. Now add the terms, and you get 12x=8. Divide by 12, you get 2/3, or 0.666666....

OpenStudy (anonymous):

Thought, and thank you both!

OpenStudy (anonymous):

Welcome